Disregarding wind resistance, the distance a body falls from rest varies directly as the square of the time it falls. If a skydiver falls 64ft in 2 seconds, how far will he fall in 10 seconds?
Answer by Mathtut(3670)

d/t=k
:
64/2=32=k
:
d/10=32
:
d=320 ft